Plz askiitians tutors help me out.... i have posted this question twice but to no avail.... its a fitjee question and only the answer is given... no detailed solution is given there... can anybody help me out  and give a detailed solution for the problem??? the questions is "Find the area bounded by y=g(x), x=1 and x=15, where g is inverse of f, and f(x)= (x^3)+3x+1..." The answer is 18...


 

Grade:12

5 Answers

Pratik Tibrewal
10 years ago
Please find the solution attached.

Thanks and Regards,
Pratik Tibrewal,
askiitians faculty,
BTech
Varun Acharya
10 years ago
i dnt understnd????
Pratik Tibrewal
10 years ago
f(x)= x^3+3x+1; g(x) is inverse of f(x). We need the area of y = g(x) b/w x=1 & x=15 It is similar to find the area of x=f^(-1) (y) b/w y =1 and y =15 Or it is similar to find area of y = f(x) b/w x = f^(-1) (1) & f^(-1) (15) And subtract this area from rectangular box 15*f^(-1) (15)-1*f^(-1) (1)-Area off(x) b/w x = f^(-1) (1) & f^(-1) (15)
Pratik Tibrewal
10 years ago
f(x)= x^3+3x+1; g(x) is inverse of f(x). We need the area of y = g(x) b/w x=1 & x=15 It is similar to find the area of x=f^(-1) (y) b/w y =1 and y =15 Or it is similar to find area of y = f(x) b/w x = f^(-1) (1) & f^(-1) (15) And subtract this area from rectangular box 15*f^(-1) (15)-1*f^(-1) (1)-Area off(x) b/w x = f^(-1) (1) & f^(-1) (15) Thanks & Regards, Pratik Tibrewal, askIITians faculty.
Varun Acharya
10 years ago
do we get the answer as 18?????